OpenStudy (anonymous):

I'm trying to negate the statement '∀ x ∈ R, ∃ y ∈ R such that x^y ∈ Z', but am wondering is it ok to say that '∀ x ∈ R, ∃ y ∈ R, such that x^y is not a member of Z' ?

OpenStudy (anonymous):

Updating with the equation formatting: \[∀ x ∈ \mathbb{R}, ∃ y ∈ \mathbb{R}\] such that \[x ^{y} ∈ \mathbb{Z}\]

OpenStudy (anonymous):

In plain English: 'for every x in the set of real numbers, there exists y in the set of real numbers such that x ^ y is in the set of integers'

OpenStudy (anonymous):

I'm not sure how to negate it is the problem

OpenStudy (anonymous):

I think I got it, I just said that for every x in the set of real numbers, there exists y in the set of real numbers such that x raised to the y is not in the set of integers